8 clocks and 5 watches cost $695 althogether. 3 clocks and 3 watches cost $300 altogether. How much more does each clock cost than each watch?
Answer by Nate(3500)

cost of watches=w
cost of clocks=c
-----------------------------
8c + 5w = 695
3c + 3w = 300
-----------------------------
combine:
24c + 15w = 2085
-15c - 15w = -1500
9c = 585
c = 65
----------------------------
3c + 3w = 300
3(65) + 3w = 300
195 + 3w = 300
3w = 105
w = 35